# 91191, Gif-sur-Yvette, France

### v1

@@ 1 Introduction and Summary
The purpose of the present work is to study a particular quantity that plays a pivotal role primarily in the Virasoro conformal bootstrap—namely, the crossing kernel of Virasoro conformal blocks. We will be interested in the crossing kernels on two particular Riemann surfaces: the torus with one marked point (modular kernel) and the sphere with four marked points (fusion kernel). The precise definitions of these kernels are given in sections 1.2 and 1.3.
The Virasoro modular and fusion kernels are mostly known through integral representations, due to the seminal works of Teschner Teschner:2003at , Ponsot-Teschner Ponsot:2000mt , and Teschner-Vartanov Teschner:2012em . As we review shortly (also in Appendix A), those formulas are build out of special meromorphic functions such as the Barnes double gamma function $\Gamma_{b}$ and the double sine function $S_{b}$ – the latter being closely related to Faddeev’s quantum dilogarithmFaddeev:1999fe – and are realized as solutions to particular shift relations (or, difference equations). They also depend in a specific way on the central charge via $c=1+6(b+b^{-1})^{2}$. On the other hand, non-integral representations for the Virasoro kernels are known only for the cases $c=1$ and $c=25$. At $c=1$, there is an interesting relation to the Painlevé VI connection constantIorgov:2013uoa ; DelMonte:2025vqv . More recently, Ribault and one of the present authors showed that fusion kernel at $c=25$ is directly related to the one at $c=1$ via a particular symmetry of the shift relations called Virasoro-Wick Rotation (see section 1.4)Ribault:2023vqs . Roughly speaking, this symmetry maps a solution of the corresponding shift relations valid at a given $c\in[25,\infty)$ to another solution valid at the corresponding $26-c\in(-\infty,1]$.
@@
$\int_{\mathcal{C}}f(z)dz=\left(\int_{\mathcal{C}}-\int_{a+\mathcal{C}}\right)\frac{f(z)}{1+\frac{f(a+z)}{f(z)}}dz.\qquad\square$ (3)
@@
We continue the rest of this section by introducing the notations and the definitions of the quantities that we will use throughout the paper. The summary of results and the organization of the paper are presented in section 1.5.
@@ 1.2 Definition of Crossing Kernels for $c\in\mathbb{C}\backslash(-\infty,1]$
The prefactor and integrand [^3] The integrand is also a function of $P_{0},P_{s},P_{t}$ but we suppress this dependence for brevity. are expressed in the terms of the Barnes double gamma function $\Gamma_{b}(x)$ and the double sine function $S_{b}(x)$ (c.f. Appendix A) as follows
@@
and analogously for $\mathrm{Re}(x)\rightarrow$-$\infty$. It can be shown that the expression (LABEL:eq:deffusionkernel) is the unique meromorphic solution to the fusion kernel shift relations for $b\in\mathbb{R}_{(0,1]}$ Eberhardt:2023mrq as consequence of the Moore-Seiberg consistency conditions.
@@ 2.1 Two novel solutions for the Modular Kernel
The definitions of $G_{m,n}(x)$ and $\widetilde{G}_{m,n}(x)$ in terms of the Barnes’ $G$ function as well as their relations with $\Gamma_{b}$ and $S_{b}$ when $b=\mathsf{b}$ are given in detail in Appendix A. In going from (LABEL:eq:defmodularkernel), (30) to (LABEL:eq:modkenrbrat), (54) we have additionally made two modifications [^9] In other words, and for the sake of clarity, we emphasize that $M_{\mathsf{b}},m_{\mathsf{b}}$ here are not exactly equal to $M_{b},m_{b}$ defined in in (30) when evaluated at $b=\mathsf{b}$.: first we multiplied and divided by a convenient $P_{s}$-dependent constant
@@
as the quantum modular determinant, and we will express various quantities in terms of it. There are of course other equivalent formulas for this determinant, and we will next present two of them, each with an interesting geometric interpretation that we will discuss shortly.
@@
The interpretation in terms of angles or lengths here requires some restriction on the values of the Liouville momenta. Indeed, we need $P_{s},P_{t},P_{0}\in\mathbb{R}$ such that $\psi_{i}\in(0,\pi)$ for the orthoscheme case, and $P_{s},P_{t},P_{0}\in\mathbb{C}$ with $\mathrm{Re}{P_{s}},\mathrm{Re}{P_{t}}=\frac{\mathbb{Z}}{2\mathsf{s}}$, and $\mathrm{Re}{P_{0}}=\frac{m+n-1}{2\mathsf{s}}+\frac{2\mathbb{Z}}{\mathsf{s}}$ for the tetrahedron case [^11] General necessary and sufficient conditions for hyperbolic orthoschemes and tetrahedra to exist are given in chokim ; murakamiyano . Among those, a distinguished one is the requirement that the determinant of (65) (for the hyperbolic orthoscheme), (67) (for the hyperbolic tetrahedron) should be strictly negative.. So far our expressions for the modular kernel have been mostly meromorphic (away from square-root branch cuts and poles) and hence there is no a priori indication that we should restrict to those values of the momenta in our analysis. Nevertheless, it is a striking and rather intriguing feature that three-polyhedra emerge in the formulas for the modular kernel at $b=\mathsf{b}$. We will return to discuss this point in Section 4.
@@ 2.4 Shift relations
For general $b^{2}$ the basic shift relation that fixes the Teschner-Vartanov fusion kernel uniquelyEberhardt:2023mrq is
@@ 3.1 Modular Kernel
$\widehat{\ell}^{(\text{m})}_{1}$ $=\widehat{\ell}^{(\text{m})}_{2}=\widehat{\ell}^{(\text{m})}_{4}=\widehat{\ell}^{(\text{m})}_{5}\equiv\pi\mathsf{s}P_{0}+\frac{i\pi}{2}(m+n-1),$ (144)
$\widehat{\ell}^{(\text{m})}_{3}$ $\equiv 4\pi\mathsf{s}P_{s},\qquad\widehat{\ell}^{(\text{m})}_{6}\equiv 4\pi\mathsf{s}P_{t}.$
The interpretation in terms of lengths here imposes $P_{s},P_{t},P_{1},P_{2},P_{3},P_{4}\in\mathbb{C}$ with $\mathrm{Im}{P_{s}},\mathrm{Im}{P_{t}}=\frac{\mathbb{Z}}{2\mathsf{s}}$ and $\mathrm{Im}{P_{0}}=\frac{1-m-n}{2\mathsf{s}}+\frac{2\mathbb{Z}}{\mathsf{s}}$. There is no indication however from our general expressions (135) that one should restrict to these particular values of the momenta. We will comment more on these interesting connections of our kernels with the three-polyhedra in Section 4.
@@ 4 Discussion
In the present work we have unravelled a novel structure of these kernels when the parameter $b^{2}$ (that controls the central charge) takes on rational values. We have explained how the original forms of the modular and fusion kernels can be seen as particular instances of a state integral, of the sort that appear in quantum topology Hirakami ; Dimofte:2009yn ; EllegaardAndersen:2011vps ; EllegaardAndersen:2013pze ; Kashaev:2012cz ; Dimofte:2014zga ; Dimofte:2015kkp . An important aspect of the crossing kernels –contrary to e.g. the Virasoro conformal blocks on which they act – is the fact that they only depend on the Liouville momenta and not on the moduli of the corresponding Riemann surface associated to conformal blocks. This makes them relatively easier to understand, especially via the shift relations that they satisfy. A proper understanding of the shift relations and the space of their solutions (or more generally, the solutions to the Moore-Seiberg consistency conditions) is pivotal, and still not a completely solved problem. Here we have explained how, if one looses the meromorphicity assumption, we can generate new solutions to those equations. What was even more striking was the fact that these new solutions (at $b^{2}\in\mathbb{Q}$) are associated with an algebraic variety – the quantum modular and fusion polynomials – and possess, via the discriminant of these polynomials, geometric-like features similar to those of three-dimensional tetrahedra. We will comment more on this particular feature next.
It is known since the work of Teschner and Vartanov Teschner:2012em that the modular and fusion kernels have an interesting “semiclassical limit”, which is defined as $b\rightarrow 0$ with the product $bP_{i}\equiv\frac{\ell_{i}}{4\pi i}$ being held fixed [^20] Recently, these semiclassical limits were nicely revisited in (Hartman:2025ula, , Appendix A). We will basically adjust their formulas in our notations.. We will be especially interested in the semiclassical limit of the integrals
@@
$\mathcal{M}_{\mathsf{b}}\left(\frac{\log{z^{(\text{m})}_{1}}}{2\pi i\mathsf{s}^{2}}\right)$ $\approx\text{exp}\left[-\frac{1}{2\pi\mathsf{s}^{2}}V_{L}\begin{pmatrix}\ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{s}\\ \ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{t}\end{pmatrix}\right$],
@@
The Virasoro modular and fusion kernels $\mathbf{M}^{(b)},\mathbf{F}^{(b)}$ are (quantum) modular forms with respect to the parameter $\tau\equiv b^{2}$.
@@ B.2 Fusion kernel
  • Good behaviour at large imaginary $u$: It is straightforward to check that the integrand in (242) behaves exactly as in (36) (with $Q\equiv Q_{m,n}=\frac{m+n}{\sqrt{mn}}$) as $u$ attains large positive/negative imaginary part. Therefore, it is safe to close the contour in the counterclockwise fashion (positive convention) at infinity, and write

### v3

@@ 1 Introduction and Summary
The purpose of the present work is to study a particular quantity that plays a pivotal role primarily in the Virasoro conformal bootstrap – namely, the crossing kernel of Virasoro conformal blocks. We will be interested in the crossing kernels on two particular Riemann surfaces: the torus with one marked point (modular kernel) and the sphere with four marked points (fusion kernel). The precise definitions of these kernels are given in sections 1.2 and 1.3.
The Virasoro modular and fusion kernels are mostly known through integral representations, due to the seminal works of Teschner Teschner:2003at , Ponsot-Teschner Ponsot:2000mt , and Teschner-Vartanov Teschner:2012em . As we will review later (also in Appendix A), those formulas are build out of special meromorphic functions such as the Barnes double gamma function $\Gamma_{b}$ and the double sine function $S_{b}$ – the latter being closely related to Faddeev’s quantum dilogarithm Faddeev:1999fe – and are realized as solutions to particular shift relations (or, difference equations). They also depend in a specific way on the central charge via $c=1+6(b+b^{-1})^{2}$. On the other hand, non-integral representations for the Virasoro kernels are known only for the cases $c=1$ and $c=25$. At $c=1$, there is an interesting relation to the Painlevé VI connection constant Iorgov:2013uoa ; DelMonte:2025vqv . More recently, Ribault and one of the present authors showed that the fusion kernel at $c=25$ is directly related to the one at $c=1$ via a particular symmetry of the shift relations called Virasoro-Wick Rotation (see section 1.4)Ribault:2023vqs . Roughly speaking, this symmetry maps a solution of the corresponding shift relations valid at a given $c\in[25,\infty)$ to another solution valid at $26-c\in(-\infty,1]$.
@@
$\int_{\mathcal{C}}f(z)dz=\left(\int_{\mathcal{C}}-\int_{a+\mathcal{C}}\right)\frac{f(z)}{1-\frac{f(a+z)}{f(z)}}dz.\qquad\square$ (3)
@@
We continue the rest of this section by introducing the notations and definitions of the quantities that we will use throughout the paper. The summary of results and the organization of the paper are presented in section 1.5.
@@ 1.2 Definition of Crossing Kernels for $c\in\mathbb{C}\backslash(-\infty,1]$
The prefactor and integrand [^3] The integrand is also a function of $P_{0},P_{s},P_{t}$ but we suppress this dependence for brevity. are expressed in terms of the Barnes double gamma function $\Gamma_{b}(x)$ and the double sine function $S_{b}(x)$ (c.f. Appendix A) as follows
@@
and analogously for $\mathrm{Re}(x)\rightarrow$-$\infty$. It can be shown that the expression (LABEL:eq:deffusionkernel) is the unique meromorphic solution to the fusion kernel shift relations for $b\in\mathbb{R}_{(0,1]}$ Eberhardt:2023mrq as a consequence of the Moore-Seiberg consistency conditions.
@@ 2.1 Two novel solutions for the Modular Kernel
The definitions of $G_{m,n}(x)$ and $\widetilde{G}_{m,n}(x)$ in terms of the Barnes’ $G$ function as well as their relations with $\Gamma_{b}$ and $S_{b}$ when $b=\mathsf{b}$ are given in detail in Appendix A. In going from (LABEL:eq:defmodularkernel), (30) to (LABEL:eq:modkenrbrat), (54) we have additionally made two modifications [^9] In other words, and for the sake of clarity, we emphasize that $M_{\mathsf{b}},m_{\mathsf{b}}$ here are not exactly equal to $M_{b},m_{b}$ defined in (30) when evaluated at $b=\mathsf{b}$.: first we multiplied and divided by a convenient $P_{s}$-dependent constant
@@
as the quantum modular determinant, and we will express various quantities in terms of it. There are of course other equivalent formulas for the discriminant, and we will next present two of them, each with an interesting geometric interpretation that we will discuss shortly.
@@
The interpretation in terms of angles or lengths here requires some restriction on the values of the Liouville momenta. Indeed, we need $P_{s},P_{t},P_{0}\in\mathbb{R}$ such that $\psi_{i}\in(0,\pi)$ for the orthoscheme case, and $P_{s},P_{t},P_{0}\in\mathbb{C}$ with $\mathrm{Re}{P_{s}},\mathrm{Re}{P_{t}}=\frac{\mathbb{Z}}{2\mathsf{s}}$, and $\mathrm{Re}{P_{0}}=\frac{m+n-1}{2\mathsf{s}}+\frac{2\mathbb{Z}}{\mathsf{s}}$ for the tetrahedron case [^11] General necessary and sufficient conditions for hyperbolic orthoschemes and tetrahedra to exist are given in chokim ; murakamiyano . Among those, a distinguished one is the requirement that the determinants (65) (for the hyperbolic orthoscheme), (67) (for the hyperbolic tetrahedron) should be strictly negative.. So far our expressions for the modular kernel have been mostly meromorphic (away from square-root branch cuts and poles) and hence there is no a priori indication that we should restrict to those values of the momenta in our analysis. Nevertheless, it is a striking and rather intriguing feature that three-polyhedra emerge in the formulas for the modular kernel at $b=\mathsf{b}$. We will return to discuss this point in Section 4.
@@ 2.4 Shift relations
For general $b^{2}$ the basic shift relation that fixes the Teschner-Vartanov fusion kernel Eberhardt:2023mrq is
@@ 3.1 Modular Kernel
$\widehat{\ell}^{(\text{m})}_{1}$ $=\widehat{\ell}^{(\text{m})}_{2}=\widehat{\ell}^{(\text{m})}_{4}=\widehat{\ell}^{(\text{m})}_{5}\equiv\widehat{\ell}^{(\text{m})}_{0}\equiv\pi\mathsf{s}P_{0}+\frac{i\pi}{2}(m+n-1),$ (144)
$\widehat{\ell}^{(\text{m})}_{3}$ $\equiv\widehat{\ell}^{(\text{m})}_{s}\equiv 4\pi\mathsf{s}P_{s},\qquad\widehat{\ell}^{(\text{m})}_{6}\equiv\widehat{\ell}^{(\text{m})}_{t}\equiv 4\pi\mathsf{s}P_{t}.$
The interpretation in terms of lengths here imposes $P_{s},P_{t},P_{0}\in\mathbb{C}$ with $\mathrm{Im}{P_{s}},\mathrm{Im}{P_{t}}=\frac{\mathbb{Z}}{2\mathsf{s}}$ and $\mathrm{Im}{P_{0}}=\frac{1-m-n}{2\mathsf{s}}+\frac{2\mathbb{Z}}{\mathsf{s}}$. There is no indication however from our general expressions (135) that one should restrict to these particular values of the momenta. We will comment more on these interesting connections of our kernels with the three-polyhedra in Section 4.
@@ 4 Discussion
In the present work we have unravelled a novel structure of these kernels when the parameter $b^{2}$ (that controls the central charge) takes on rational values. We have explained how the original forms of the modular and fusion kernels can be seen as particular instances of a state integral, of the sort that appear in quantum topology hikami ; Dimofte:2009yn ; EllegaardAndersen:2011vps ; EllegaardAndersen:2013pze ; Kashaev:2012cz ; Dimofte:2014zga ; Dimofte:2015kkp . An important aspect of the crossing kernels –contrary to e.g. the Virasoro conformal blocks on which they act – is the fact that they only depend on the Liouville momenta and not on the moduli of the corresponding Riemann surface associated to conformal blocks. This makes them relatively easier to understand, especially via the shift relations that they satisfy. A proper understanding of the shift relations and the space of their solutions (or more generally, the solutions to the Moore-Seiberg consistency conditions) is pivotal, and still not a completely solved problem. Here we have explained how, if one looses the meromorphicity assumption, we can generate new solutions to those equations. What was even more striking was the fact that these new solutions (at $b^{2}\in\mathbb{Q}$) are associated with an algebraic variety – the quantum modular and fusion polynomials – and possess, via the discriminant of these polynomials, geometric-like features similar to those of three-dimensional tetrahedra. We will comment more on this particular feature next.
It is known since the work of Teschner and Vartanov Teschner:2012em that the modular and fusion kernels have an interesting “semiclassical limit”, which is defined as $b\rightarrow 0$ with the product $bP_{i}\equiv\frac{\ell_{i}}{4\pi i}$ being held fixed [^20] Recently, these semiclassical limits were revisited in Hartman:2025ula ; Liu:2025tzv . Here we will basically follow (Hartman:2025ula, , Appendix A).. We will be especially interested in the semiclassical limit of the integrals
@@
$\mathcal{M}_{\mathsf{b}}\left(\frac{\log{z^{(\text{m})}_{1}}}{2\pi i\mathsf{s}^{2}}\right)$ $\approx\text{exp}\left[-\frac{1}{2\pi\mathsf{s}^{2}}V_{L}\begin{pmatrix}\ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{s}\\ \ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{0}&\ell^{(\text{m})}_{t}\end{pmatrix}\right$], (175)
@@
The Virasoro modular and fusion kernels $\mathbf{M}^{(b)},\mathbf{F}^{(b)}$ are associated to (quantum) modular forms with respect to the parameter $\tau\equiv b^{2}$.
@@ B.2 Fusion kernel
  • Good behaviour at large imaginary $u$: It is straightforward to check that the integrand in (243) behaves exactly as in (36) (with $Q\equiv\mathsf{Q}=\frac{m+n}{\sqrt{mn}}$) as $u$ attains large positive/negative imaginary part. Therefore, it is safe to close the contour in the counterclockwise fashion (positive convention) at infinity, and write